Find the distance between (–4, 0) and (–5, –3)
thanks

Answer by checkley75(3666)

the x distance is (5-4)=1
the y distance is (3-0)=3
therefore the distance is found by the pythagoreon theorem
a^2+b^2=c^2 a=x distance, b=y distance & c=hypotenuse
1^2+3^2=c^2
1+9=c^2
c^2=10
c=sqrt10
c=3.16 answer for the hypotenuse ofr distance between the given points.